Derive the relation between impulse and momentum.
1
Expert's answer
2020-09-10T14:40:24-0400

Using the Newton's second law:


"\vec{F}=m\frac{d\vec{v}}{dt}"

We can rewrite it as:


"\vec{F}dt=md\vec{v}""\int\vec{F}dt=\int md\vec{v}""\vec{F}\Delta t=m\Delta \vec{v}"

Which is by definition of impulse and momentum:


"\vec{J}=\Delta \vec{p}"

So, the impulse equals to the change of momenta.